### v1

@@ 3 Conclusion and Outlook
In this work, we image the cross-DoF structure in spatial-polarization hyperentangled biphotons produced from SPDC using a data-driven camera capable of time-stamping single photons. We first demonstrate the rapid characterization of hyperentanglement with a cumulative measurement time of only 17 minutes. In particular, we verify transverse spatial entanglement through a violation of the Einstein-Podolsky-Rosen criterion with a minimum conditional uncertainty product of (0.110.05) and certify the entanglement dimensionality to be about 148 in the spatial DoF using the method discussed in Refs. [44, 45]. By performing spatially-resolved polarization state tomography, we confirm the generation of an entire class of near-maximally polarization-entangled states with an average concurrence of 0.83030.0004 in the entire spatial profile of the SPDC field, thereby certifying a spatial-polarization hyperentanglement with an estimated total dimensionality of 251. We then present the cross-influence between spatial and polarization DoF. By producing the first complete spatial map of the polarization state generated from SPDC, we illustrate the dependence of polarization entanglement on the transverse momenta of the down-converted photons in the entire spatial profile of the SPDC field. Using a weakly focused pump beam, we, for the first time, experimentally characterize the cross-influence between the polarization DoF of the down-converted photons and the spatial DoF of the pump beam.
@@
4 Methods
4.1 Detailed experimental setup
An ultraviolet continuous-wave laser emits the pump beam with a center wavelength of 405 nm, bandwidth of 2 nm, and power of 20 mW. The pump beam is made polarized at using a polarizing beam splitter (PBS) and a half-wave plate (HWP). A 5-mm temporal compensator (TC) quartz crystal introduces a time delay between the horizontal (H-) and vertical (V-) components of the pump beam, which pre-compensates for the temporal walk-off that the two orthogonal polarization components are expected to subsequently gather inside the BBO double-crystal [60]. The BBO crystals are each 0.5-mm-thick and identically cut for type-I phase-matching, with their optics axes oriented perpendicularly to one another [26]. A -polarized photon from the pump beam can then be down-converted in either crystal with equal probability to produce a pair of H-polarized or V-polarized photons. Factoring in the multiple momentum-correlated spatial modes of the down-converted photons, the output state is given by Eq. (1). In the experiment, we orient the double-crystal for near-collinear phase matching.
The down-converted photons are separated from the pump beam using a dichroic mirror (DM), and then probabilistically split into two arms using a non-polarizing beam splitter (BS). We ignore the cases wherein both the down-converted photons exit the same port of the BS, and denote photons that are reflected as “signal” and those that are transmitted as “idler”. The signal and idler photons then propagate through a quarter-wave plate (QWP) and a HWP before they are either transmitted or reflected by a PBS. The photons finally pass through a bandpass filter with a center wavelength of 800 nm and full width at half maximum (FWHM) bandwidth of 40 nm before being detected by the Tpx3Cam. To characterize the entanglement in the spatial DoF, we implement two imaging schemes for realizing measurements in the position (near-field) and momentum (far-field) bases. Using the combination of mm, mm, and mm lenses, the far-field of the crystal is imaged on the Tpx3Cam, thereby enabling measurements in the momentum basis. On the other hand, using mm images the crystal output face (near-field) onto the Tpx3Cam sensor, thereby allowing for measurements in the position basis.
The imaging sensor in Tpx3Cam comprises a pixel array with a pixel pitch of 55 m. The pixels in Tpx3Cam are data-driven and individually trigger the registration of photon incidence events when the signal amplitude exceeds a predefined threshold. The camera can be single photon sensitive with the addition of an image intensifier (Photonis Cricket) and have a single photon temporal resolution of 2 ns [31]. Time-stamping for individual photons can potentially allow for more versatility in data acquisition and analysis compared to frame-based imaging devices.
4.2 Data acquisition and processing with Tpx3Cam
To characterize entanglement in the spatial DoF, we set the waveplates to project both the signal and idler photons into the V-polarization basis and acquire data in the near-field and far-field for 1 minute each using the Tpx3Cam. To characterize entanglement in the polarization DoF, we additionally acquire data for 15 different combinations of orientations of HWPs and QWPs in the far-field, each for 1 minute. Combined with the VV basis data acquired in characterizing the far-field spatial correlation, these polarization measurements allow us to reconstruct the polarization density matrices corresponding to all pairs of momentum-anticorrelated pixels through quantum state tomography and characterize polarization entanglement. The total required data acquisition time for characterizing a spatial-polarization hyperentangled state is thus minutes.
During data acquisition, each incident photon could hit a cluster of pixels due to being amplified by the intensifier. To correct this, we apply a centroiding algorithm that identifies the amplitude-weighed center in each cluster as the true pixel coordinate. We use the time stamp of each centroided pixel as a reference to correct the time walk within the cluster. After centroiding and time walk correction, we apply a two-pointer technique to the sorted time stamps of signal and idler photons and identify events detected within a 10-ns time window as photon coincidences [32]. Although the time walk correction algorithm can reduce the temporal resolution to 8 ns [57, 61], it should not significantly affect the efficiency of the subsequent coincidence counting with our choice of longer time window.
4.3 Estimation of fidelity lower bound
Following the approaches described in [46], we certify the dimensionality of entanglement in the spatial DoF using correlations in two MUBs [45, 44]. In Sec. 2.2, we have defined the two MUBs to be the discrete momentum basis and the discrete position basis .
The fidelity of the experimentally measured state to a maximally entangled state is defined as:
where
Since for an arbitrary state with a Schmidt number of , the fidelity satisfies
the entanglement dimensionality of a state with must be at least . As we will show in the following, correlations in the two MUBs will allow us to obtain a lower bound for the fidelity via
The coincidence counts measured in the discrete momentum basis allow us to calculate using
Supplementing the coincidence counts measured in the discrete position basis allows us to calculate , the lower bound of , via
where
As a result, we calculate a lower bound of the fidelity , which certifies the entanglement dimensionality to be 148.
5 Data availability
All relevant data are available from the corresponding author upon reasonable request.
6 Code availability
All relevant codes are available from the corresponding author upon reasonable request.
7 Acknowledgements
@@
The authors thank fruitful discussion with X.Gao, A.D’Errico, B. Braverman, S. Karan, and M.Krenn. C.L. acknowledges P. Svihra for helpful advice on data processing.
8 Author contribution
C.L. conceived the research idea; C.L. and I.S. conducted the experiments with help from Y.Z., D.E., A.N. and E.K. in setting up the Tpx3Cam system for data acquisition; C.L., G.K., Y.Z., J.U., D.E. and A.N. analyzed the data; C.L. and G.K. wrote the first draft of the manuscript and all authors subsequently contributed to the writing; D.E., A.N., E.K. and R.B. supervised the project.

### v2

@@ 4 Conclusion and Outlook
In this work, we image the cross-DoF structure in spatial-polarization hyperentangled biphotons produced from SPDC using a data-driven camera capable of time-stamping single photons. We first demonstrate the rapid characterization of hyperentanglement with a cumulative measurement time of only 17 minutes. In particular, we certify the entanglement dimensionality to be about 148 in the spatial DoF using the method discussed in Refs. [48, 49]. By performing spatially-resolved polarization state tomography, we confirm the generation of an entire class of near-maximally polarization-entangled states with an average concurrence of 0.83030.0004 in the entire spatial profile of the SPDC field, thereby certifying a spatial-polarization hyperentanglement with an estimated total dimensionality of 251. We then present the cross-influence between spatial and polarization DoF. By producing the first complete spatial map of the polarization state generated from SPDC, we illustrate the dependence of polarization entanglement on the transverse momenta of the down-converted photons in the entire spatial profile of the SPDC field. Using a weakly focused pump beam, we, for the first time, experimentally characterize the cross-influence between the polarization DoF of the down-converted photons and the spatial DoF of the pump beam.
@@
Funding National Science Foundation (2244031); U.S. Department of Energy (76295); Natural Sciences and Engineering Research Council of Canada (RGPIN/2017-06880); Canada First Research Excellence Fund (072623).
@@
Acknowledgement The authors thank X. Gao, A. D’Errico, B. Braverman, S. Karan, and M. Krenn for fruitful discussions. C. Li acknowledges P. Svihra for helpful advice on data processing.
Data availability All relevant data are available from the corresponding author upon reasonable request.
Disclosures The authors declare no conflict of interest.